OpenStudy (anonymous):

If A,B are the roots of px^2+px+r=0, form an equation whose roots are (2A+1/A), (AB+1/B) . . Answer= prx^2+q(2r+p)x+(2r-p)^2+2q^2

OpenStudy (anonymous):

Formula for forming the equation x^2-(sum of roots)x+(product of roots)=0

OpenStudy (anonymous):

I am stuck at product of roots

OpenStudy (anonymous):

sum=-q/p product=r/p

OpenStudy (anonymous):

Sum of given roots: (2A+1/A)+(2B+1/B)

OpenStudy (anonymous):

\[(2A+2B) +(\frac{ 1 }{ A }+\frac{ 1 }{ B })\]

OpenStudy (anonymous):

I got the sum of roots which was \[\frac{ -q(2r+p) }{ pr }\]

OpenStudy (anonymous):

having problem in product of roots

OpenStudy (anonymous):

(2A+1/A)*(2B+1/B)

OpenStudy (anonymous):

anyone there?

OpenStudy (anonymous):

its been an hour since I asked the question and got no answer yet :/

hartnn (hartnn):

(2A+1/A)*(2B+1/B) what u get after multiplying them out ?

OpenStudy (anonymous):

oh finally someone came to my Q

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

I am showing step by step what I did

OpenStudy (anonymous):

4AB + 2A/B +2B/A +1/AB

hartnn (hartnn):

so, u only have problem with 2A/B +2B/A part, right ?

OpenStudy (anonymous):

YUP

OpenStudy (anonymous):

problem in simplification of product of roots

OpenStudy (anonymous):

(4AB + 1/AB) + (2A/B + 2B/A)

hartnn (hartnn):

so, 2 [A^2+B^2]/AB so you just need, A^2+B^2, right ?

OpenStudy (anonymous):

(4AB^2+1 / AB ) + ( 2A^2+2B^2/AB)

hartnn (hartnn):

note that you already know AB =....

hartnn (hartnn):

so, you know (4AB + 1/AB)

OpenStudy (anonymous):

yes and A+B too

hartnn (hartnn):

so, finally , you just need A62+B62, right ?

hartnn (hartnn):

A^2+B^2

OpenStudy (anonymous):

\[\frac{ 4(\frac{ r }{ p })^2+1 }{ \frac{ r }{ p } }\]

OpenStudy (anonymous):

yes it will expand by formula..right?

OpenStudy (anonymous):

(4A^2+4B^2)-8AB

hartnn (hartnn):

huh ? what ?

hartnn (hartnn):

(4AB + 1/AB) + (2A/B + 2B/A) = 4 (r/p) + p/r +(2A/B + 2B/A) right ?

OpenStudy (anonymous):

\[2A+\frac{ 1 }{ A } * 2B+\frac{ 1 }{ B }\]

hartnn (hartnn):

don't miss the brackets..... are u clear till here ? (4AB + 1/AB) + (2A/B + 2B/A) = 4 (r/p) + p/r +(2A/B + 2B/A)

hartnn (hartnn):

AB = r/p so, 1/AB = p/r

OpenStudy (anonymous):

OK

hartnn (hartnn):

now you need (2A/B + 2B/A) = 2 [A^2+B^2]/AB right ?

OpenStudy (anonymous):

OKAY

hartnn (hartnn):

now use the fact that A^2+B^2 = (A+B)^2 - 2AB you know both A+B and AB

OpenStudy (anonymous):

A^2+B^2 WILL EXPAND BY FORMULA??

OpenStudy (anonymous):

OK .OK

hartnn (hartnn):

try it, if you don't get it now, then tell me.

OpenStudy (anonymous):

I SIMPLIFIED IT

OpenStudy (anonymous):

and got it till \[\frac{ 4r^2+p^2+2q^2-2rp }{ pr }\]

OpenStudy (raden):

px^2+px+r=0 where is the q ? i didnt see it :P

hartnn (hartnn):

i assumed px^2+qx+r=0

OpenStudy (anonymous):

sorry @RadEn TYPING PROBLEM

OpenStudy (anonymous):

I DID IT (2r-p)^2+2q^2/pr

hartnn (hartnn):

*applauses*